OpenStudy (anonymous):

What are the exact solutions of x2 = 5x + 2?

OpenStudy (mathmate):

You can transpose the equation to the form: \(ax^2+bx+c=0\) and use the quadratic formula.

OpenStudy (mathmate):

Recall the quadratic formula is: \(\huge x=\frac{-b\pm\sqrt{b^2-4ac}}{2a} \)